# Enabling Approximate Joint Sampling in Diffusion LMs

### v1

@@ Abstract
In autoregressive language models, each token is sampled by conditioning on all the past tokens; the overall string has thus been sampled from the correct underlying joint distribution represented by the model. In contrast, masked diffusion language models generate text by unmasking tokens out of order and potentially in parallel. Generating an overall string sampled from the correct underlying joint distribution would (again) require exactly one token unmasking in every full-model forward pass. The more tokens unmasked in parallel, the further away the string is from the true joint; this can be seen in the resulting drop in accuracy (but, increase in speed). In this paper we devise a way to approximately sample multiple tokens from the joint distribution in a single full-model forward pass; we do so by developing a new lightweight single-layer “sampler” on top of an existing large diffusion LM. One forward pass of the full model can now be followed by multiple forward passes of only this sampler layer, to yield multiple unmasked tokens. Our sampler is trained to mimic exact joint sampling from the (frozen) full model. We show the effectiveness of our approximate joint sampling for both pretrained-only (Dream-7B-Base) and instruction-tuned (Dream-7B-Instruct) models on language modeling and math & coding tasks. When four tokens are unmasked for each full-model denoising step, our sampling algorithm achieves a MAUVE score of 0.87 (vs marginal baseline of 0.31) with respect to the true joint distribution.
@@ 1 Introduction
Masked diffusion language models Sahoo et al. (2024); Austin et al. (2021); Lou et al. (2023) involve generating text strings by starting from an all-masked sequence of tokens, and then iteratively replacing the masked tokens with tokens from the vocabulary, with each “denoising” forward pass unmasking one or a few tokens. As opposed to auto-regressive models which generate tokens left to right and one token in each forward pass, in masked diffusion models tokens can be potentially unmasked in any order and also potentially multiple tokens can be unmasked in parallel.
The higher the number of tokens unmasked in parallel after a single denoising forward pass, the faster and cheaper the overall generation Sahoo et al. (2024). However, the quality of the resulting generation is also generally seen to be worse. State-of-the-art large masked diffusion models (7B parameters) are advertised Ye et al. (2025); Nie et al. (2025) to achieve competitive downstream task performance, when compared to their auto-regressive counterparts; however, this parity is only achieved via slow one-token-per-denoise-step generation from the masked diffusion model. On decreasing the number of denoising steps, there is a sharp decrease in their performance. For example, Dream-7B-Instruct model’s accuracy drops from 85% accuracy to 77% when generating two tokens instead of one per denoising step Israel et al. (2025).
@@
    We show the utility of our method ADJUST on both unconditional generation (Sec 4.2) and downstream evaluation (Sec 4.3). Across models, sampling parameters and evaluation settings, ADJUST yields consistent gains compared to naive parallel sampling. For example, when considering four tokens generated per time step, ADJUST increases GSM8k accuracy by 16 points and MAUVE score by 0.5 points. In the absence of domain-specific prompts, a generically trained (assuming input prompt to be null) ADJUST, is able to generalize and perform well on downstream tasks as well.
@@ 2 Related Work
Discrete diffusion language models (LMs) Sahoo et al. (2024); Shi et al. (2024); Lou et al. (2023) serve as a promising alternative to auto-regressive (AR) LMs. Large diffusion LMs Ye et al. (2025); Nie et al. (2025) competing with the AR counterparts are based on an absorbing-state Markov process, which noises tokens into a “masked” state. Masked diffusion language models can also be considered as mask denoisers  Zheng et al. (2024). Discrete Copula Diffusion Liu et al. (2024b) uses both AR and diffusion models in tandem to get final logits which have the best of both worlds (future conditioned and joint probabilities). One line of work focuses on how to select the next token to unmask and propose trainable planners for achieving the best downstream performance Peng et al. (2025); Liu et al. (2024c). Apart from diffusion language models, we also mention other work closely related to this paradigm such as any-order auto-regressive models Shih et al. (2022), among others  Sahoo et al. (2025); Chao et al. (2025)
Efficient decoding for language models
Prior work has proposed techniques for speeding up diffusion language models by introducing KV cache and adaptive decoding of confident tokens Wu et al. (2025); Israel et al. (2025). These works leverage the confidence score of the diffusion model or an off-the-shelf AR model to decide how many tokens to commit often speeding up inference. Speculative decoding is a popular paradigm for provably accelerating auto-regressive models Leviathan et al. (2023). Prior work on this has developed drafter techniques which heavily inspire our method Li et al. (2025; 2024); Liu et al. (2024a); Zhang et al. (2024). Prior work also introduces unrolling the draft model during training (to better align test and train) which we use in our work Li et al. (2025); Liu et al. (2024a); Zhang et al. (2024). Recent work also proposes speculative decoding for any-order models Guo & Ermon (2025).
Figure 2: In this figure we report runtime (tokens produced per second) with the corresponding MAUVE and GSM8k scores. We vary the number of tokens sampled for a single diffusion denoising step to vary the token throughput (from one to four tokens per step, each corresponding to a point on the curve). Increasing the number of tokens sampled per step leads to a decrease in both MAUVE and GSM8k performance, as expected. We show that our joint sampler leads to only slight reduction in throughput compared to the naive parallel sampling. Importantly, for a given target throughput, sampling using our joint approximation outperforms using naive parallel sampling. For example, for parallel decoding of four tokens in each diffusion pass, our joint approximation is only 20-25% slower than parallel decoding while being 16 % more accurate on GSM8K and 0.5 points higher on MAUVE.
@@ 3.1 Problem Setup
In this paper we only consider base models that are masked diffusion models with “unmask and commit” inference; that is, the tokens to be generated are initially started out as M, and these locations are iteratively filled in with other tokens from the vocabulary via forward passes of the model. Unconditional generation would thus involve a series of strings $\phi\rightarrow{\bm{x}}_{n_{1}}\rightarrow{\bm{x}}_{n_{2}}\rightarrow\ldots\rightarrow{\bm{x}}_{L}$, where $\phi={\bm{x}}_{0}$ is the all-mask sequence and ${\bm{x}}_{L}$ is the completely unmasked string. In this paper we will build a joint sampler for an existing, frozen base diffusion model which we will refer to by the pair $(f,{\bm{W}})$. Here $f:{\mathcal{V}}^{L}\to\mathbb{R}^{d\times L}$111we assume that the time information is implicit in the masked input sequence Zheng et al. (2024) refers to the map from an input string of $L$ tokens $\bm{x}$} to an output of $L$ embeddings, with $f^{i}({\bm{x}})$ denoting the output embedding at the $i^{th}$ position. ${\bm{W}}\in\mathbb{R}^{{\mathcal{V}}\times d}$ is the language model head that maps an embedding to logits. Note that diffusion models only output per-token marginals; that is for each position $i$ a single forward pass yields a probability distribution $p^{i}(\cdot|{\bm{x}})=\mathrm{softmax}[{\bm{W}}f^{i}({\bm{x}})]$ over the vocabulary $\mathcal{V}$. If multiple tokens are unmasked in the same forward pass, this is done by independently generating each of them from their respective marginals.
@@
$p_{parallel}\left(x^{\sigma(1)},\ldots,x^{\sigma(K)}\right)~=~\prod_{k=1}^{K}\,p^{\sigma(k)}\left(x^{\sigma(k)}\,|\,\phi\right)~=~\prod_{k=1}^{K}\,p_{*}\left(x^{\sigma(k)}\,|\,\phi\right)$
@@
Figure 3: In this figure we illustrate our method ADJUST. The sub-figure on the left shows naive parallel sampling, while the sub-figure on the right shows our method ADJUST. In this example, each denoising step generates three tokens. The underlying diffusion model has support on only two distinct sentences, “The cat sat on the mat” and “The dog ran in the yard”. Sampling tokens in parallel (left), generates an incoherent sentence (w.r.t. the underlying distribution of diffusion model). Diffusion model’s support does not contain a sentence which mentions a “cat”, “on the yard”. On the other hand, our method ADJUST (right) conditions each token sample on the previously sampled tokens through a light-weight network, shown in the figure as $g$. Note that we utilize the same number of forward passes of the diffusion model $f$ as the naive parallel sampler (once every three generated tokens).
3.2 Our approximate joint sampler
@@
The ADJUST model is represented as the pair $(g,{\bm{W}})$, similar to how the base diffusion model is $(f,{\bm{W}})$. The function $g:\mathbb{R}^{d\times L}\times{\mathcal{V}}^{L}\to\mathbb{R}^{d\times L}$ takes two inputs: ${\bm{h}}\in\mathbb{R}^{d\times L}$, which is a set of “most current” embeddings (elaborated on later), as well as the currently unmasked string (say $\tilde{{\bm{x}}}$). $g$ represents the output embeddings of ADJUST. Let $q$ denote the output marginal probabilities of ADJUST. Then we can denote the probabilities as :
[…]
where $\bm{h}$} and $\tilde{{\bm{x}}}$ are the most current embeddings and token strings respectively. We draw the reader’s attention to the parallels between $q$ and $p$, where the only difference between the two is the inclusion of the extra embeddings $\bm{h}$} as input for $g$.
Let us now walk through how ADJUST  sequentially generates tokens; let $\sigma(1),\ldots,\sigma(K)$ be the set of positions it will unmask. The first token $\tilde{x}^{\sigma(1)}$ can be unmasked directly using the base model forward pass $p^{\sigma(1)}(\cdot\mid{\bm{x}})$ itself. After sampling the first token, the most current token string gets updated to ${\bm{x}}\oplus\tilde{x}^{\sigma(1)}$, while our most current embedding is $f({\bm{x}})$. From the second token onward our method ADJUST kicks-in. The most current embeddings are first updated (with the most current string) using ADJUST as :
${\bm{h}}_{2}=g\left(f({\bm{x}}),{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\right)$
and then the second token $\tilde{x}^{\sigma(2)}$ is sampled using the logits ${\bm{W}}{\bm{h}}_{2}$. This means that
$p_{\text{ADJUST}}(\tilde{x}^{\sigma(1)},\tilde{x}^{\sigma(2)}\mid{\bm{x}})~=~p^{\sigma(1)}\left(\tilde{x}^{\sigma(1)}\mid{\bm{x}}\right)\>q^{\sigma(2)}\left(\tilde{x}^{\sigma(2)}\mid f({\bm{x}}),{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\right)$
Now for sampling of the third token, our most current embeddings are ${\bm{h}}_{2}$ and most current token string is ${\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\tilde{x}^{\sigma(2)}$. These are used as input during the next iteration of $g$. Hence, letting ${\bm{h}}_{1}=f({\bm{x}})$, we can define the recursion for output embeddings of $g$ as
${\bm{h}}_{k}=g\left({\bm{h}}_{k-1}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$
@@
$p_{\text{ADJUST}}(\tilde{x}^{\sigma(1)},\ldots,\tilde{x}^{\sigma(K)}\mid{\bm{x}})$~
$=$~ $p^{\sigma(1)}\left(\tilde{x}^{\sigma(1)}\,|\,{\bm{x}}\right)\prod_{k=2}^{K}\,q^{\sigma(k)}\left(\tilde{x}^{\sigma(k)}\mid{\bm{h}}_{k-1}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$
Note that for sampling $K$ tokens, we do a single forward pass through the base diffusion model, for computing ${\bm{h}}_{1}=f({\bm{x}})$ and K1K-1 forward passes through our ADJUST $g$ for computing all subsequent ${\bm{h}}_{k}$. We present our joint sampling algorithm in Alg 1.
@@
3:  ${\bm{h}}_{1}\leftarrow f({\bm{x}})$ $\triangleright$ Compute output embeddings of the base model
4:  for $k=1$ to $K$ do
5:   $i\leftarrow{\mathrm{TOP}}\{{\bm{W}}{{\bm{h}}}^{j}_{k}:j\in\mathcal{M}({\bm{x}})\}$ $\triangleright$ Select index to unmask
@@
7:   ${\bm{x}}\leftarrow{\bm{x}}\oplus\tilde{x}^{i}$ $\triangleright$ Get updated token string
8:   ${{\bm{h}}}_{k+1}\leftarrow g({{\bm{h}}}_{k},{\bm{x}})$ $\triangleright$ Update output embeddings using ADJUST
@@
We clarify that our method is NOT a speculative decoding Leviathan et al. (2023) method. Speculative decoding refers to a “guess and verify” paradigm for accelerating inference through auto-regressive (AR) language models. It is specifically useful for AR models as the verification (NLL computation) of a guessed string (a single forward pass through the AR model) is cheaper than auto-regressive generation (K forward passes). Since diffusion models do not provide an efficient way to verify tokens which is cheaper than generating the tokens themselves, speculative decoding techniques are not applicable for diffusion models. We highlight two other differences between ADJUST and speculative decoding. Firstly, as shown above, we always accept all the tokens generated by ADJUST (there is no verification). Recall that diffusion models, through a single forward pass, give us a marginal distribution (prior belief) for each masked position. Our formulation considers conditioning this prior belief on the sampled tokens. This is in contrast to drafters used for AR models in speculative decoding, where the goal is to speculate future tokens from scratch (no prior belief is given by the base model).

### v2

@@ Abstract
In autoregressive language models, each token is sampled by conditioning on all the past tokens; the overall string has thus been sampled from the correct underlying joint distribution represented by the model. In contrast, masked diffusion language models generate text by unmasking tokens out of order and potentially in parallel. Generating an overall string sampled from the correct underlying joint distribution would (again) require exactly one token unmasking in every full-model forward pass. The more tokens unmasked in parallel, the further away the string is from the true joint; this can be seen in the resulting drop in accuracy (but, increase in speed). In this paper we devise a way to approximately sample multiple tokens from the joint distribution in a single full-model forward pass; we do so by developing a new lightweight single-layer “sampler” on top of an existing large diffusion LM. One forward pass of the full model can now be followed by multiple forward passes of only this sampler layer, to yield multiple unmasked tokens. Our sampler is trained to mimic exact joint sampling from the (frozen) full model. We show the effectiveness of our approximate joint sampling for both pretrained-only (Dream-7B-Base, Llada-7B-Base) and instruction-tuned (Dream-7B-Instruct, Dream-7B-Coder) models on downstream tasks (GSM8k, MATH, MBPP, HEval) and on unconditional language modeling. When eight tokens are unmasked for each full-model denoising step, our sampling algorithm achieves a MAUVE score of 0.84 (vs marginal baseline of 0.19) with respect to the true joint distribution.
Machine Learning, ICML
@@ 1 Introduction
Masked diffusion language models (Sahoo et al., 2024; Austin et al., 2021; Lou et al., 2023) involve generating text strings by starting from an all-masked sequence of tokens, and then iteratively replacing the masked tokens with tokens from the vocabulary, with each “denoising” forward pass unmasking one or a few tokens. As opposed to auto-regressive models which generate tokens left to right and one token in each forward pass, in masked diffusion models tokens can be potentially unmasked in any order and also potentially multiple tokens can be unmasked in parallel.
The higher the number of tokens unmasked in parallel after a single denoising forward pass, the faster and cheaper the overall generation (Sahoo et al., 2024). However, the quality of the resulting generation is also generally seen to be worse. State-of-the-art large masked diffusion models (7B parameters) are advertised (Ye et al., 2025; Nie et al., 2025; Song et al., 2025; Khanna et al., 2025) to achieve competitive downstream task performance, when compared to their auto-regressive counterparts; however, this parity is only achieved via slow one-token-per-denoise-step generation from the masked diffusion model. On decreasing the number of denoising steps, there is a sharp decrease in their performance. For example, Dream-7B-Instruct model’s accuracy drops from 85% accuracy to 77% when generating two tokens instead of one per denoising step (Israel et al., 2025).
@@
Figure 2: In this figure we report runtime (tokens produced per second) with the corresponding MAUVE and GSM8k scores for Dream-Base-7B and Dream-Instruct-7B respectively. We vary the number of tokens sampled for a single diffusion denoising step to vary the token throughput (from one to four tokens per step, each corresponding to a point on the curve). Increasing the number of tokens sampled per step leads to a decrease in both MAUVE and GSM8k performance, as expected. We show that our joint sampler leads to only slight reduction in throughput compared to the naive parallel sampling. Importantly, for a given target throughput, sampling using our joint approximation outperforms using naive parallel sampling. For example, for parallel decoding of four tokens in each diffusion pass, our joint approximation is only 20-25% slower than parallel decoding while being 16 percentage points more accurate on GSM8K and 0.5 points higher on MAUVE.
@@
    We show the utility of our method ADJUST on both unconditional generation (Sec 4.2) and downstream evaluation (Sec 4.3). Across models, sampling parameters and evaluation settings, ADJUST yields consistent gains compared to naive parallel sampling. For example, when considering four tokens generated per time step, ADJUST increases GSM8k accuracy by 16 percentage points and MAUVE score by 0.5 points. In the absence of domain-specific prompts, a generically trained (assuming input prompt to be null) ADJUST, is able to generalize and perform well on downstream tasks as well.
@@ 2 Related Work
Discrete diffusion language models (LMs) (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023) are a promising alternative to auto-regressive (AR) LMs. Large diffusion LMs (Ye et al., 2025; Nie et al., 2025) competing with the AR counterparts are based on an absorbing-state Markov process, which noises tokens into a “masked” state.  Zheng et al. (2024) show that masked diffusion language models can also be considered as mask denoisers. In order to further the capabilities of diffusion LMs Liu et al. (2024b); Israel et al. (2025) propose to use both AR and diffusion models in tandem to get final logits which have the best of both auto-regressive and diffusion worlds (future conditioned and joint probabilities). Other works try to resolve the pressing challenges of diffusion language models.  Rout et al. (2025) show that latent conditioning on important ”anchor” tokens leads to much better performance.
Parallel Token Generation
Prior work has proposed techniques for speeding up auto-regressive language models by decoding multiple tokens for each forward pass of the auto-regressive (AR) model (Samragh et al., 2025; Li et al., 2025, 2024; Liu et al., 2024a; Zhang et al., 2024). These techniques are used in a speculative decoding paradigm  (Leviathan et al., 2023) where the generated tokens are verified before accepting them.  Israel et al. (2025); Guo and Ermon (2025) propose methods very similar to speculative decoding for non-AR with the goal of improving token latency. These works leverage the confidence score of the diffusion model or an off-the-shelf AR model to decide how many tokens to commit often speeding up inference.  Campbell et al. (2025); Arriola et al. (2025) train specialized models which are cheaper to infer. At a high level they consider a ”encoder” backbone along with an ”decoder” model. Our work instead proposes parallel decoding as joint sampling from the underlying distribution of diffusion model.
@@ 3 Method
Figure 3: In this figure we illustrate our method ADJUST. The sub-figure on the left shows naive parallel sampling, while the sub-figure on the right shows our method ADJUST. In this example, each denoising step generates three tokens. For this illustration, assume that the diffusion model $f$ has support on only two distinct sentences, “The cat sat on the mat” and “The dog ran in the yard”. Sampling tokens in parallel (left), generates an incoherent sentence with respect to the underlying distribution of diffusion model. On the other hand, our method ADJUST (right) conditions each token sample on the previously sampled tokens through a light-weight network, shown in the figure as $g$. Note that we utilize the same number of forward passes of the diffusion model $f$ as the naive parallel sampler (once every three generated tokens).
@@ 3.1 Problem Setup
In this paper we only consider base models that are masked diffusion models with “unmask and commit” inference; that is, the tokens to be generated are initially started out as M, and these locations are iteratively filled in with other tokens from the vocabulary via forward passes of the model. Unconditional generation would thus involve a series of strings $\phi\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{2}\rightarrow\ldots\rightarrow{\bm{x}}_{T}$, where $\phi={\bm{x}}_{0}$ is the all-mask sequence and ${\bm{x}}_{T}$ is the completely unmasked string. In this paper we will build a joint sampler for an existing, frozen base diffusion model which we will refer to by the pair $(f,{\bm{W}})$. Here $f:{\mathcal{V}}^{L}\to\mathbb{R}^{d\times L}$ 111we assume that the time information is implicit in the masked input sequence (Zheng et al., 2024) refers to the map from an input string of $L$ tokens $\bm{x}$} to an output of $L$ embeddings, with $f^{i}({\bm{x}})$ denoting the output embedding at the $i^{th}$ position. ${\bm{W}}\in\mathbb{R}^{{\mathcal{V}}\times d}$ is the language model head that maps an embedding to logits. Note that diffusion models only output per-token marginals; that is for each position $i$ a single forward pass yields a probability distribution $p^{i}(\cdot|{\bm{x}})=\mathrm{softmax}[{\bm{W}}f^{i}({\bm{x}})]$ over the vocabulary $\mathcal{V}$. If multiple tokens are unmasked in the same forward pass, this is done by independently generating each of them from their respective marginals.
@@
$p_{parallel}\left(x^{\sigma(1)},\ldots,x^{\sigma(K)}\right)$ $=$ $\prod_{k=1}^{K}\,p^{\sigma(k)}\left(x^{\sigma(k)}\,|\,\phi\right)$
@@
3.2 Our approximate joint sampler : ADJUST
@@
The ADJUST model is represented as the pair $(g,{\bm{W}})$, similar to how the base diffusion model is $(f,{\bm{W}})$. The function $g:\mathbb{R}^{d\times L}\times{\mathcal{V}}^{L}\to\mathbb{R}^{d\times L}$ takes two inputs: ${\bm{h}}\in\mathbb{R}^{d\times L}$, which is a set of “most current” embeddings (elaborated on later), as well as the currently unmasked string (say $\bm{x}$}). Then for each position i, ADJUST yields a probability distribution :
$q^{i}\left(\cdot\mid{\bm{h}},{{\bm{x}}}\right)=\text{softmax}\big[{\bm{W}}g^{i}({\bm{h}},{{\bm{x}}})\big]$
where $\bm{h}$} and $\bm{x}$}} are the most current embeddings and token strings respectively. We draw the reader’s attention to the parallels between $q$ and $p$, where the only difference between the two is the inclusion of the extra embeddings $\bm{h}$} as input for $g$.
Sampling with ADJUST Let $\sigma(1),\ldots,\sigma(K)$ be the set of positions it will unmask. The first token $\tilde{x}^{\sigma(1)}$ can be unmasked directly using the base model forward pass $p^{\sigma(1)}(\cdot\mid{\bm{x}})$ itself. After sampling the first token, the most current token string gets updated to ${\bm{x}}\oplus\tilde{x}^{\sigma(1)}$, while our most current embedding is $f({\bm{x}})$. From the second token onward our method ADJUST kicks-in. The most current embeddings are first updated (with the most current string) using ADJUST as : ${\bm{h}}_{1}=g\left(f({\bm{x}}),{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\right)$ and then the second token $\tilde{x}^{\sigma(2)}$ is sampled from ${\bm{W}}{\bm{h}}^{\sigma(2)}_{1}$ (where ${\bm{h}}^{j}$ denotes the $j$th index of $\bm{h}$}). Hence, letting ${\bm{h}}_{0}=f({\bm{x}})$, we can define the recursion for output embeddings of $g$ as
${\bm{h}}_{k+1}=g\left({\bm{h}}_{k}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k+1)}\right)$
@@
$p_{\text{ADJUST}}(\tilde{x}^{\sigma(1)},\ldots,\tilde{x}^{\sigma(K)}\mid{\bm{x}})~=~p^{\sigma(1)}\left(\tilde{x}^{\sigma(1)}\,|\,{\bm{x}}\right)\times$
$\prod_{k=2}^{K}\,q^{\sigma(k)}\left(\tilde{x}^{\sigma(k)}\mid{\bm{h}}_{k-2}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$
Note that for sampling $K$ tokens, we do a single forward pass through the base diffusion model, for computing ${\bm{h}}_{1}=f({\bm{x}})$ and K1K-1 forward passes through our ADJUST $g$ for computing all subsequent ${\bm{h}}_{k}$. We present our joint sampling algorithm in Alg 1. Note that the only difference between naive and ADJUST sampling is updating the output embeddings (highlighted in different color).
@@
3:  ${\bm{h}}_{0}\leftarrow f({\bm{x}})$
4:  for $k=0$ to K1K-1 do
5:   $i\leftarrow{\mathrm{TOP}}\{{\bm{W}}{{\bm{h}}}^{j}_{k}:j\in\mathcal{M}({\bm{x}})\}$
@@
7:   ${\bm{x}}\leftarrow{\bm{x}}\oplus\tilde{x}^{i}$
8:   ${{\bm{h}}}_{k+1}\leftarrow g({{\bm{h}}}_{k},{\bm{x}})$
@@
We clarify that our method is NOT a speculative decoding (Leviathan et al., 2023) method. Speculative decoding refers to a “guess and verify” paradigm for accelerating inference through auto-regressive (AR) language models. It is specifically useful for AR models as the verification (NLL computation) of a guessed string (a single forward pass through the AR model) is cheaper than auto-regressive generation (K forward passes). Since diffusion models do not provide an efficient way to verify tokens which is cheaper than generating the tokens themselves, speculative decoding techniques are not applicable for diffusion models. We highlight two other differences between ADJUST and speculative decoding. Firstly, as shown above, we always accept all the tokens generated by ADJUST (there is no verification). Recall that diffusion models, through a single forward pass, give us a marginal distribution (prior belief) for each masked position. Our formulation considers conditioning this prior belief on the sampled tokens. This is in contrast to drafters used for AR models in speculative decoding, where the goal is to speculate future tokens from scratch (no prior belief is given by the base model).